# Effective-medium approach to the resonance distribution of wave scattering in a random point field

### v2

@@ Abstract
In a previous paper [Phys. Rev. A 105, 042205 (2022)], the distribution of resonance poles in the complex plane of the wavenumber $k$ associated to the multiple scattering of a quantum particle in a random point field was numerically discovered. This distribution presented two distinctive structures: a set of peaks at small $k$ when the wavelength is larger than the interscatterer distance, and a band almost parallel to the real axis at larger $k$. In this paper, a detailed theoretical study based on wave transport theory is proposed to explain the origin of these structures and to predict their distribution in the complex $k$ plane. First, it is shown that the peaks at small $k$ can be understood using an effective wave equation for the average wavefunction over the disorder. Then, that the band at large $k$ can be described by the Bethe-Salpeter equation for the square modulus of the wavefunction, which is derived from the diagrammatic method. This study is supported by careful comparisons with numerical simulations. The largest simulations revealed the presence of filaments alike quantum scars for the wavefunction in the bulk of the disordered medium.
Resonance poles; Disordered medium; Effective medium; Quantum transport; Multiple scattering; Diffusion; Random Lorentz gas; Random point field; Point scatterers; Foldy-Lax model; Bethe-Salpeter equation
@@ I Introduction
The propagation of waves in disordered media is a subject of crucial importance in many research areas in physics such as acoustics [1, 2, 3, 4, 5], electromagnetism [6, 7, 8, 9, 10], and matter waves [11, 12, 13, 14, 15, 16, 17, 18]. These phenomena are described by wave equations with random or statistically correlated spatial heterogeneities in their propagation parameters such as the index of refraction or the potential energy function [19, 20, 21, 22, 23, 24, 25, 26, 27]. Examples are given by multiple scattering of waves in a random field of scatterers such as atoms or molecules in gases, liquids, or amorphous solids. Disorder is known to cause special effects on wave propagation, depending on the ratio of the wavelength over the mean distance between the scatterers. Pioneering works on multiple scattering by Foldy [28] and Lax [29, *Lax1952] have shown that a disordered medium may be considered under some conditions as an effective medium characterized by spatially averaged propagation parameters such as a mean refractive index. Later on, Anderson discovered that waves may be localized as the result of multiple scattering in low-dimensional disordered media [31, 32]. Otherwise, in semiclassical regimes, wave multiple scattering leads to transport ruled by the diffusion equation or the Boltzmann kinetic equation [33, 34, 35].
@@
The purpose of the present paper is to show that these structures in the resonance distribution can be quantitatively explained with accuracy using an extension of the effective-medium theory from real to complex values for the wavenumber. To our knowledge, such a generalization of the effective-medium theory does not seem to be known in the literature. The effective-medium wave equation, also referred to as the Dyson equation, allows us to understand the formation of the low-energy structures previously observed in the resonance distribution. Furthermore, we can show that the diffusion approximation to the transport equation, known in this research area as the Bethe-Salpeter equation [19, 20, 21, 22, 27, 49, 50, 51, 52, *Doicu2019a2, *Doicu2019a3], can explain the mean spatial profile of the wavefunction density obtained after averaging over many random configurations of the system, as well as key features of the resonance distribution at high energies. Furthermore, we also evaluate the coherence length, resulting from the process of decoherence due to the randomness of the configuration of the scatterers in the disordered medium.
This paper is organized as follows. The Foldy-Lax model used to describe the propagation of the quantum particle in a gaseous detector is presented in Sec. II.1. An example of wavefunction obtained numerically in the Foldy-Lax model for a disk-shaped random point field is shown in Sec. II.2. The scattering resonances in the complex plane of the wavenumber are defined in Sec. II.3. In Sec. III, an effective-medium theory is derived from averaging over the random configurations of the atoms to predict the distribution of peaks in the resonance density at low wavenumber. This theory is limited to a single scattering event and must be supplanted, in the large wavenumber regime, by a transport theory for the multiple scattering. This is the topic of Sec. IV which contains several steps: First the presentation of semiclassical transport equations in Secs. IV.1, IV.2, and IV.3, and then comparisons with numerical simulations in Secs. IV.4 and IV.5. Finally, conclusions are drawn in Sec. V. This paper also contains several appendices with detailed derivations, consistency checks, and additional explanations regarding the equations of the main text.
@@
where $\Gamma(z)$ denotes the gamma function [55]. The unit $d$-ball itself will be denoted by $\mathcal{B}_{d}$, and its external border by $\mathcal{S}_{d}$.
@@ II.1 The Foldy-Lax model
In order to describe the multiple collisions of the quantum particle in a disordered medium, we consider a simple but powerful model originally introduced by Foldy and Lax [28, 29, *Lax1952, 27]. In this model, the wavefunction of the particle undergoes elastic collisions without loss of energy in a random configuration of fixed point scatterers representing the atoms or the molecules of a gaseous detector. This model is very convenient especially because it can be efficiently solved numerically for large disordered systems without resorting to huge discretization lattices. In this model, the wavefunction $\psi(\boldsymbol{\mathrm{r}})$ obeys the stationary Schrödinger equation
@@
$n(\boldsymbol{\mathrm{r}})=\begin{cases}N/V&\text{if}~{}\boldsymbol{\mathrm{r}}\in\mathcal{B}_{d}(R)\>,\\ 0&\text{otherwise}\>,\end{cases}$ (6)
where $V=V_{d}R^{d}$ is the volume of the random point field using the notation (1). In all numerical simulations, we will use the mean interparticle distance
@@
In all numerical simulations, unless otherwise mentioned, we choose to maximize the scattering amplitude (12) by imposing $\cot\delta(k)=0$:
@@
II.2 Single-configuration wavefunction
An example of the square modulus of the wavefunction for a large random configuration of $N=10^{4}$ scatterers is shown in Fig. 1(a), using the incident wave
@@ II.2 Single-configuration wavefunction
Figure 1: Square modulus of the wavefunction computed with the program MSModel [56] in a 2D disk-shaped random point field for a single configuration of $N=10^{4}$ scatterers with $k=10\,\varsigma^{-1}$. The point source is at the center ($\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$) and the mean free path is $\ell_{\rm s}=2.5\,\varsigma$. The circle depicts the medium boundary. (a) 4-megapixel overview. Total computation time is about 80 minutes on an Intel Core i7 processor with four threads. (b) Enlargement of the central source point. The scatterers are depicted by small black dots. (c) Enlargement of a presumed quantum scar.
The point-like scatterers are depicted as dots in Fig. 1. As seen in this figure, the density of the single-configuration wavefunction strongly fluctuates at the scale of the wavelength. In particular, the scale of the angular fluctuations of $\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}$ out of the medium ($\left\|\boldsymbol{\mathrm{r}}\right\|\gg R$) is predicted by partial-wave theory to be
$\Delta\theta=\frac{\pi}{kR}\>,$ (23)
as given in Eq. (137) of Ref. [46]. In the case of Fig. 1(a), this gives a fluctuation scale of the order of $\Delta\theta\simeq 0.3^{\circ}$.
The figure also shows that, within the medium, the wavefunction presents disordered meandering nodal lines, which are separated on average by a distance of the order of half the wavelength, i.e., about $\pi/k$. In addition, we observe filament-shaped enhancements of the magnitude of the wavefunction density. These structures look alike the quantum scars observed in the wavefunctions of classically chaotic systems, where the wavefunction magnitude may be enhanced by constructive interferences in the vicinity of unstable classical periodic orbits [61, *Heller2001]. There is no such orbit in the present model, since the cross section of points is classically equal to zero. However, the wave nature of the quantum particle allows a nonzero cross section, which is at most of the order of $\lambda^{d-1}$, as we have seen in Eq. (16). The cross section is thus nonzero for $\lambda\neq 0$, so that the role of unstable periodic orbits may here be played by closed paths joining several point scatterers, which might explain the similarity between the structures of nevertheless different kinds of quantum systems. It should be noted that the random point field contains [^1] There are $\textstyle\binom{N}{p}=\frac{N!}{p!(N-p)}$ ways to pick up $p$ points out of $N$. This count is then multiplied by the number $\tfrac{1}{2}(p-1)!$ of closed cycles in a complete graph of $p$ vertices.
$\mathcal{C}_{{\rm cy},p}=\begin{cases}\frac{N!}{2p(N-p)!}&\text{if}~{}p\geq 3\>,\\ \tfrac{1}{2}N(N-1)&\text{if}~{}p=2\>,\end{cases}$ (24)
closed paths involving $p$ points. The total number of closed paths thus reads
$\mathcal{C}_{{\rm cy}}=\sum_{p=2}^{N}\mathcal{C}_{{\rm cy},p}\xrightarrow{N\rightarrow\infty}\frac{(N-1)!}{2}\mathop{}\!\mathrm{e}^{1}\>.$ (25)
The number (25) is so large in the limit $N\rightarrow\infty$, that we may expect such scarring structures to occur in many possible places. It is probably the first time a phenomenon of quantum scarring is observed for such a large disordered system.
The computation of such wavefunctions can be repeated for many different random configurations of the scatterers to perform their statistical average.
II.3 Scattering resonances
This paper primarily aims at studying the distribution of resonances which are formally defined as the singular points, or poles, of the resolvent $\hat{Q}(k)$ of the complete multiple-scattering problem (2)–(3). This resolvent operator is defined by
@@ II.3 Scattering resonances
Equation (29) has infinitely many roots in the complex plane of $k$, and the density $\varrho^{(2)}({k}_{\rm r},{k}_{\rm i})$ of these roots increases with the number $N$ of scatterers. The average of this two-dimensional distribution over the random configurations of the disorder can be obtained with the resonance potential method introduced in Ref. [47]. This average resonance distribution displays the structures seen, for instance, in Fig. 6 of Ref. [47].
III Effective-medium approach
In this section, we develop a theory for the low-energy peaks in the complex resonance density observed in our previous paper [47]. These peaks appear at low energy, that is in a regime of large wavelength compared to the typical interscatterer distance ($k\varsigma\ll 1$). In this regime, we expect that the wave is no longer able to resolve the spatial heterogeneities of the medium, and thus perceives a uniform effective medium [26]. A possible approach to derive an effective-medium wave equation is to average directly the wavefunction over the random configurations of the scatterers, hence leading to a Dyson-type wave equation [19, 20, 21, 26, 25, 52, *Doicu2019a2, *Doicu2019a3]. As we will see, this equation successfully predicts the locations of the resonance peaks at low energy, providing a quantitative understanding of the numerical observations in our previous paper [47].
III.1 Effective-medium wave equation
The idea of the effective-medium approach is to define the amplitude function
$A(\boldsymbol{\mathrm{r}})=\sum_{i=1}^{N}a_{i}\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{x}}_{i})\>,$ (30)
which will hopefully tend to a continuous function upon averaging over the random configurations of the scatterers. Let us express Eq. (18) in terms of $A(\boldsymbol{\mathrm{r}})$. Multiplying both sides by $\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{x}}_{i})$ and summing over $i$ gives us
$\begin{split}A(\boldsymbol{\mathrm{r}})&=F(k)\left(\sum_{i=1}^{N}\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{x}}_{i})\phi(\boldsymbol{\mathrm{r}})\right.\\ &\left.+\sum_{i=1}^{N}\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{x}}_{i})\sum_{j(\neq i)}^{N}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{x}}_{j})a_{j}\right)\>.\end{split}$ (31)
The sum over $i$ can be averaged over the random configurations. Since we assume that the point field is uniform with density $n$ given by Eq. (5), Eq. (31) averages to
$\left\langle A(\boldsymbol{\mathrm{r}})\right\rangle=n(\boldsymbol{\mathrm{r}})F(k)\left(\phi(\boldsymbol{\mathrm{r}})+\left\langle\sum_{j(\neq i)}^{N}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{x}}_{j})a_{j}\right\rangle\right)\>.$ (32)
Note that, to get Eq. (32), we neglected the statistical correlation between $a_{j}$ and $\boldsymbol{\mathrm{x}}_{i}$ in the last term of Eq. (31). In order to close the equation for $\langle A(\boldsymbol{\mathrm{r}})\rangle$, we need to achieve another approximation on the last term of Eq. (32). We will assume that, under the configurational average, the sum over $\boldsymbol{\mathrm{x}}_{j}$ can be replaced by an integral
$\left\langle\sum_{j(\neq i)}^{N}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{x}}_{j})a_{j}\right\rangle\simeq\int_{\mathbb{R}^{d}}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\left\langle A(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>.$ (33)
The integral in the right-hand side of Eq. (33) ignores the exclusion of the $i$-th scatterer in the sum of the left-hand side. To some extent, this difference could be compensated by a factor $(N-1)/N$ in the right-hand side, but we deliberately omit it here because the number of scatterers is supposed to be large enough to make this factor tend to 1 [28, 29]. Finally, substituting Eq. (33) into Eq. (32) leads to
$\begin{split}\left\langle A(\boldsymbol{\mathrm{r}})\right\rangle&=n(\boldsymbol{\mathrm{r}})F(k)\\ &\times\left(\phi(\boldsymbol{\mathrm{r}})+\int_{\mathbb{R}^{d}}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\left\langle A(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\right)\>.\end{split}$ (34)
This is a self-consistent equation for $\langle A(\boldsymbol{\mathrm{r}})\rangle$. However, we want an equation for the average wavefunction itself. We return for this purpose to Eq. (17) and we average it over the configurations. The result is
$\left\langle\psi(\boldsymbol{\mathrm{r}})\right\rangle=\phi(\boldsymbol{\mathrm{r}})+\int_{\mathbb{R}^{d}}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\left\langle A(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>.$ (35)
Note that, in contrast to Eq. (34), the average of Eq. (17) is not subject to any approximation. This is due to the fact that Eq. (17) refers to individual scatterer positions $\boldsymbol{\mathrm{x}}_{i}$, whereas Eq. (31) involves pairs of scatterers $(\boldsymbol{\mathrm{x}}_{i},\boldsymbol{\mathrm{x}}_{j})$. When Eq. (35) is compared to Eq. (34), one notices they are directly proportional to each other. One finds
$\left\langle A(\boldsymbol{\mathrm{r}})\right\rangle=n(\boldsymbol{\mathrm{r}})F(k)\left\langle\psi(\boldsymbol{\mathrm{r}})\right\rangle\>.$ (36)
Therefore, Eqs. (34) and (36) imply that the average wavefunction obeys the self-consistent equation
$\left\langle\psi(\boldsymbol{\mathrm{r}})\right\rangle=\phi(\boldsymbol{\mathrm{r}})+\int_{\mathbb{R}^{d}}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})n(\boldsymbol{\mathrm{r}}^{\prime})F(k)\left\langle\psi(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>.$ (37)
It is worth noting that this equation closely resembles the general Lippmann-Schwinger equation but with $n(\boldsymbol{\mathrm{r}})F(k)$ in the role of the potential. Therefore, this means that the average wavefunction $\langle\psi(\boldsymbol{\mathrm{r}})\rangle$ is effectively governed by the same Schrödinger equation (2) as the actual wavefunction $\psi(\boldsymbol{\mathrm{r}})$, but with $U(\boldsymbol{\mathrm{r}})$ replaced by $n(\boldsymbol{\mathrm{r}})F(k)$. From Eq. (37), one deduces the Dyson-type effective-medium equation [19, 20, 21, 26, 25, 52, *Doicu2019a2, *Doicu2019a3]
$\left[\nabla^{2}_{\boldsymbol{\mathrm{r}}}+k^{2}-n(\boldsymbol{\mathrm{r}})F(k)\right]\left\langle\psi(\boldsymbol{\mathrm{r}})\right\rangle=0\>.$ (38)
However, the important difference between $U(\boldsymbol{\mathrm{r}})$ and $n(\boldsymbol{\mathrm{r}})F(k)$ is that the latter is complex while the former is real. As a consequence, Eq. (38) possesses exponentially evanescent solutions. Of course, this does not mean that the particle is absorbed in the medium, since the probability is strictly conserved in the model. In fact, this exponential evanescence of the wave is due to the destructive interferences that results from the disorder. The non-vanishing part of the average wavefunction $\langle\psi(\boldsymbol{\mathrm{r}})\rangle$ can be interpreted as the coherent wave, because it keeps in phase with the incident wave $\phi(\boldsymbol{\mathrm{r}})$ [29, 63, 26, 27, 52, *Doicu2019a2, *Doicu2019a3].
III.2 Disorder-averaged Green function
In wave transport theory, a more important quantity than the average wavefunction $\langle\psi(\boldsymbol{\mathrm{r}})\rangle$ is the full Green function averaged over the realizations of the disorder. As shown in Appendix A using the diagrammatic method, the average Green function $\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$ obeys the effective-medium equation (38) but with a point source at $\boldsymbol{\mathrm{r}}^{\prime}$ in the right-hand side:
@@ III.2 Disorder-averaged Green function
$\left[\nabla^{2}_{\boldsymbol{\mathrm{r}}}+k^{2}-n(\boldsymbol{\mathrm{r}})F(k)\right]\left\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\right\rangle=\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime})\>.$ (39)

### v3

@@ Abstract
In a previous paper [Phys. Rev. A 105, 042205 (2022)], the distribution of resonance poles in the complex plane of the wavenumber $k$ associated to the multiple scattering of a quantum particle in a random point field was numerically discovered. This distribution presented two distinctive structures: a set of peaks at small $k$ when the wavelength is larger than the interscatterer distance, and a band almost parallel to the real axis at larger $k$. In this paper, a theoretical study based on wave transport theory is proposed to explain the origin of these structures and to predict their distribution in the complex $k$ plane. First, it is shown that the peaks at small $k$ can be understood using the effective wave equation for the average wavefunction over the disorder. Then, that the band at large $k$ can be described by the Bethe-Salpeter equation for the square modulus of the wavefunction. This study is supported by careful comparisons with numerical simulations.
Resonance poles; Disordered medium; Effective medium; Quantum transport; Multiple scattering; Diffusion; Random Lorentz gas; Random point field; Point scatterers; Foldy-Lax model; Dyson equation; Bethe-Salpeter equation
@@ I Introduction
The propagation of waves in disordered media is a subject of crucial importance in many research areas in physics such as acoustics [1, 2, 3, 4, 5], electromagnetism [6, 7, 8, 9, 10], and matter waves [11, 12, 13, 14, 15, 16, 17, 18]. These phenomena are described by wave equations with random or statistically correlated spatial heterogeneities in their propagation parameters such as the index of refraction or the potential energy function [19, 20, 21, 22, 23, 24, 25, 26, 27]. Examples are given by multiple scattering of waves in a random field of scatterers such as atoms or molecules in gases, liquids, or amorphous solids. Disorder is known to cause special effects on wave propagation, depending on the ratio of the wavelength over the scattering mean free path. Pioneering works on multiple scattering by Foldy [28] and Lax [29, *Lax1952] showed that a disordered medium may be considered under some conditions as an effective medium characterized by spatially averaged propagation parameters such as an effective refractive index. Later on, Anderson discovered that waves may be localized as the result of multiple scattering in disordered media [31, 32]. Otherwise, in semiclassical regimes, wave multiple scattering leads to transport ruled by the diffusion equation or the Boltzmann kinetic equation [33, 34, 35].
@@
The purpose of the present paper is to show that these structures in the resonance distribution can be quantitatively explained with accuracy using the effective-medium theory at complex values of the wavenumber. The effective-medium wave equation, also referred to as the Dyson equation [19, 20, 21, 22, 26, 27, 49, *Doicu2019a2, *Doicu2019a3], allows us to understand the formation of the low-energy structures previously observed in the resonance distribution. Furthermore, we can show that the diffusion approximation to the transport equation, known in this research area as the Bethe-Salpeter equation [19, 20, 21, 22, 25, 26, 27, 52, 53, 54, 49, *Doicu2019a2, *Doicu2019a3], can explain key features of the resonance distribution at high energies.
This paper is organized as follows. The Foldy-Lax model used to describe the propagation of the quantum particle in a gaseous detector is presented in Sec. II.1. The scattering resonances in the complex plane of the wavenumber are defined in Sec. II.2. The disorder-averaged Green function is introduced in Sec. II.3. This function is at the heart of the semiclassical multiple-scattering theory for the square modulus of the Green function developed in Sec. II.4. Both disorder-averaging approaches are used to explain the structures observed in the resonance distribution. The predictions of the theory are then compared in Sec. III with the exact resonance distributions computed numerically. Finally, conclusions are drawn in Sec. IV.
@@
where $\Gamma(z)$ denotes the gamma function [55].
@@ II.1 The Foldy-Lax model
In order to describe the multiple collisions of the quantum particle in a disordered medium, we consider a simple but powerful model originally introduced by Foldy and Lax [28, 29, *Lax1952, 25]. In this model, the wavefunction of the particle undergoes elastic collisions without loss of energy in a random configuration of fixed point scatterers representing the atoms or the molecules of a gaseous detector. This model is very convenient especially because it can be efficiently solved numerically for large disordered systems without resorting to huge discretization lattices. In this model, the wavefunction $\psi(\boldsymbol{\mathrm{r}})$ obeys the stationary Schrödinger equation
@@
$n(\boldsymbol{\mathrm{r}})=\begin{cases}N/V&\text{if}~{}\boldsymbol{\mathrm{r}}\in\mathcal{V}\>,\\ 0&\text{otherwise}\>,\end{cases}$ (6)
where $V$ is the volume of $\mathcal{V}$. In all numerical simulations, we will use the mean interparticle distance
@@
The phase shift in the point scattering amplitude (12) intrinsically depends on the microscopic model used to describe the scattering over the individual scatterers. When the radius of the individual scatterers is much smaller than the wavelength, the scattering amplitude universally behaves as [46]
$F_{\rm hs}(k)=-\frac{I(k,\alpha)}{I(k,0)G^{+}(k,\alpha)}\>,$ (15)
where $I(k,r)=-\operatorname{Im}G^{+}(k,r)$ and $\alpha$ is known as the scattering length. This quantity must be smaller than the wavelength for this model to be valid ($\alpha k\ll 1$). Beside this, the point cross section (14) can be maximized by imposing $\cot\delta(k)=0$ in Eq. (12). This provides the parameter-free scattering model:
@@
An example of the square modulus of the wavefunction for a large random configuration of $N=10^{3}$ scatterers located in a two-dimensional (2D) disk-shaped region is shown in Fig. 1(a) using the incident wave
@@
Figure 1: (a) Square modulus of the wavefunction, $\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}$, computed from Eqs. (18), (20), and (23) using the program MSModel [56] in a 2D disk-shaped random point field of $N=10^{3}$ scatterers with $k=10\,\varsigma^{-1}$. The point scattering model is Eq. (16) and the mean free path is thus $\ell_{\rm s}=2.5\,\varsigma$. The circle depicts the medium boundary. (b) Same settings as panel (a) but averaged over 6464 6464 disorder configurations.
Figures similar to Fig. 1 exist in the literature for other systems [61]. The average of the square modulus of the wave function over several random configurations of the scatterers is shown in panel 1(b). This panel illustrates in particular that disorder averaging restores the spherical symmetry of the problem due to the uniformity of the scatterer density (6).
II.2 Scattering resonances
This paper aims at studying the distribution of resonances which are formally defined as the singular points, or poles, of the resolvent $\hat{Q}(k)$ of the complete multiple-scattering problem (2)–(3). This resolvent operator is defined by
@@ II.2 Scattering resonances
Equation (27) has infinitely many roots in the complex plane of $k$, and the amount of these roots increases with the number $N$ of scatterers. We thus define the complex resonance density as [47]
@@
$\varrho^{(2)}({k}_{\rm r},{k}_{\rm i})=\frac{1}{N}\sum_{p=1}^{\infty}\left\langle\delta^{(2)}(k-k_{p})\right\rangle\>,$ (28)
where the sum is over the resonance poles and $\delta^{(2)}(k-k_{p})$ is the two-dimensional Dirac delta over the complex plane of $k$. The two-dimensional distribution (28) can be obtained numerically with the resonance potential method [47] which consists in the evaluation of
$\varrho^{(2)}({k}_{\rm r},{k}_{\rm i})=\frac{1}{2\pi N}\left(\frac{\partial^{2}}{\partial{k}_{\rm r}^{2}}+\frac{\partial^{2}}{\partial{k}_{\rm i}^{2}}\right)\left\langle\ln\left|\det\mathsf{M}(k)\right|\right\rangle\>.$ (29)
This average resonance distribution displays the structures seen, for instance, in Fig. 6 of Ref. [47], and is the focus of interest in this paper.
II.3 Disorder-averaged Green function
In this section, we develop a theory for the low-energy peaks in the complex resonance density observed in our previous paper [47]. These peaks appear at low energy, that is in a regime of large wavelength compared to the typical interscatterer distance ($k\varsigma\ll 1$). In this regime, we expect that the wave is no longer able to resolve the spatial heterogeneities of the medium, and thus perceives a uniform effective medium [26]. A possible approach to derive an effective-medium wave equation is to average directly the wavefunction over the random configurations of the scatterers, hence leading to a Dyson-type wave equation [19, 20, 21, 26, 27, 49, *Doicu2019a2, *Doicu2019a3]. As we will see, this equation successfully predicts the locations of the resonance peaks at low energy, providing a quantitative understanding of the numerical observations in our previous paper [47].
In wave transport theory, an important quantity is the full Green function averaged over the realizations of the disorder. It can be shown that the average Green function $\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$ obeys the Dyson-type effective-medium equation [28, 29, *Lax1952, 19, 20, 21, 26, 27, 49, *Doicu2019a2, *Doicu2019a3, 62, 63]
$\left[\nabla^{2}_{\boldsymbol{\mathrm{r}}}+k^{2}-n(\boldsymbol{\mathrm{r}})F(k)\right]\left\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\right\rangle=\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime})\>,$ (30)